To meet its commitments under the Copenhagen Agreement, the U.S. has to implement an excise tax on ozone depleters such as the chemical CFC-12. Originally, the market for CFC-12 is:
Demand Qd = 18.4 – 0.5P Supply Qs = 10 + 2.5P
After a 60 cent (0.6) excise tax is imposed, the supply curve shifts to:
Post-tax Supply Qst = 8.5 + 2.5P
a) What is the equilibrium price P before the tax is imposed?
b) What is the equilibrium price P after the tax is imposed?
c) Who bears the bigger burden of the tax – buyers or sellers?
d) What is the deadweight loss of this tax?
1
Expert's answer
2016-10-21T12:45:09-0400
Qd = 18.4 – 0.5P, Qs = 10 + 2.5P, t = $0.6, Qst = 8.5 + 2.5P
a) The equilibrium price P before the tax is imposed is:
Qd = Qs,
18.4 – 0.5P = 10 + 2.5P,
P = 8.4/3 = $2.8,
Q = 10 + 2.5*2.8 = 17 units.
b) The equilibrium price P after the tax is imposed is:
Qd = Qst,
18.4 – 0.5P = 8.5 + 2.5P,
P = 9.9/3 = $3.3,
Q = 8.5 + 2.5*3.3 = 16.75 units.
c) As the increase in price is 3.3 - 2.8 = $0.5 and the tax imposed is $0.6, then buyers bears the bigger burden of the tax or $0.5 from every unit sold.
d) The deadweight loss of this tax is DWL = 0.5*(3.3 - 2.8)*(17 - 16.75) = $0.0625.